OpenStudy (instagrammodel):

Evaluate each expression using a calculator.. 1024 1 -- 5 I dont know how to do this please help <33

OpenStudy (phi):

do you mean \[ 1024^\frac{1}{5} \]?

OpenStudy (instagrammodel):

Yes! haha, sorry :)

OpenStudy (instagrammodel):

@phi

OpenStudy (phi):

you can do this in your head (if you know how) but with a calculator you type in 1024 then the ^ sign (for "exponent") then either ( 1/5) or 0.2 then =

OpenStudy (instagrammodel):

oh wow! I feel so stupid! Wow.. Thank you so much. so sorry for the stupid question!

OpenStudy (phi):

but it is well-known (to those who know it) 2^10 = 1024 and (2^10)^(1/5) = 2^(10/5) = 2^2 = 2*2= 4